# Fine-Tune an SLM or Prompt an LLM? The Case of Generating Low-Code Workflows

### v1

@@ Abstract
Fine-Tune an SLM or Prompt an LLM?
The Case of Generating Low-Code Workflows
Orlando Marquez Ayala  Patrice Bechard  Emily Chen  Maggie Baird  Jingfei Chen ServiceNow {orlando.marquez, patrice.bechard, emily.chen, maggie.baird, jingfei.chen}@servicenow.com
1 Introduction
Generative AI is pushing the boundaries of AI-based software. As Large Language Models (LLMs) become more capable, they have become an integral part of the software stack. Although prompting state-of-the-art models such as GPT-4o Hurst et al. (2024), Gemini 2.0 Flash Gemini et al. (2023), or Claude 3.5 Sonnet Anthropic (2024) generally suffice for traditional tasks like question-answering or summarization Fu et al. (2024), it is not always the case on domain-specific tasks Trad and Chehab (2024).
While building Flow Generation [^1] https://www.servicenow.com/docs/csh?topicname=flowgeneration.html&version=latest, an application that generates low-code workflows based on textual user requirements, we explored whether fine-tuning a language model was necessary to achieve the desired software quality. To reduce infrastructure needs, we only considered Small Language Models (SLMs), having less than 15 billion parameters.
Figure 1: Task consists of generating a complete workflow from a user requirement.
@@ 1 Introduction
The simplest approach to generate low-code workflows is to use prompt engineering on an off-the-shelf LLM. This would be satisfactory if the steps were high-level descriptions and the logic elements were simple. But to obtain reasonable results in this task, we had to decompose it into sub-tasks Wies et al. (2023) and use Retrieval-Augmented Generation (RAG) Gao et al. (2023) at various points in the pipeline. We found out that, while prompting an LLM gives reasonable results, we can do better by fine-tuning an SLM.
Our approach was to create a small yet representative training dataset for the sub-tasks comprising Flow Generation and fine-tune Mistral-Nemo-12B-Base Mistral (2024). For evaluation, we labeled about a thousand workflows coming from ten domains (around 100 per domain). We also created a special dataset by inviting expert enterprise users to interact with the tool and collecting data generated from their usage. We compare our approach against an instruction fine-tuned SLM as the baseline, as well as against a variety of closed and open-source LLMs, showing the benefits of fine-tuning. As our task is domain-specific, we devised a metric called Flow Similarity (FlowSim), a version of tree edit distance that represents workflows as trees.
@@
2 Related Work
Prompting McCann et al. (2018); Brown et al. (2020) has emerged as the predominant method for using LLMs, significantly reducing the effort required compared to fine-tuning Sanh et al. (2021); Wei et al. (2021). However, for domain-specific applications requiring extensive prior knowledge, practitioners often encounter limitations such as constrained context windows Peng et al. (2023), which hinder the model’s ability to capture the full scope of the task. To address this, alternative approaches such as RAG Lewis et al. (2020) or domain-specific fine-tuning Gururangan et al. (2020) are employed to incorporate relevant knowledge. In various industry-specific settings, specialized models have demonstrated superior performance over state-of-the-art general-domain LLMs Wu et al. (2023); Tu et al. (2024); Colombo et al. (2024). We add to this body of work by showing that a fine-tuned SLM performs better than prompting LLMs in low-code workflow generation.
@@ 2 Related Work
Evaluation methodologies vary across tasks, with each task having its own quantitative metric for model quality but also its own qualitative error analysis approach Dou et al. (2024); Vilar et al. (2006); Gauthier-melancon et al. (2022); Bolya et al. (2020). In this work, we outline an error analysis approach for a structured output task and leverage it to gain deeper insights into model failure modes.
3 Methodology
@@ 3.1 Task Decomposition
Figure 2: Sample JSON representation of the trigger and first component steps for the workflow in Figure 1.
3.2 Retrieval-Augmented Generation
It has been shown that RAG is necessary to reduce hallucinations when generating structured outputs Ayala and Bechard (2024); Bassamzadeh and Methani (2024). In our pipeline, RAG is used to suggest the following to the LLM:
@@ 3.3 Pipeline
Fine-tuning the SLM and the retriever was done using standard approaches: supervised instruction fine-tuning and contrastive loss training.
Figure 3: System architecture with UI, AI, and data layers.
4 Experiments
4.1 Datasets
@@
Table 1: Number of examples for training (Internal and Synthetic) and evaluation (TEST and OOD).
4.2 Models
For SLMs, we considered several models such as StarCoderBase-7B Li et al. (2023) and Llama-3.1-8B Grattafiori et al. (2024), but we chose Mistral-Nemo-12B-Base Mistral (2024) because this model architecture is well-supported and optimized in our enterprise system. However, we obtained similar results with other SLMs as the key ingredient for our fine-tuning is the training data.
As the baseline, we used a model of the same architecture and size that has been instruction fine-tuned but not on data from our workflow domain: Mistral-Nemo-12B-Instruct Mistral (2024). This allows us to validate that domain-specific data is helpful. As for LLMs, we try to cover the current landscape by comparing against:
@@ 4.2 Models
  • Open-source: Llama-3.3-70B-Instruct Grattafiori et al. (2024)
  • Reasoning: o3-mini OpenAI (2025) with medium reasoning
@@
Table 2: Number of tokens in templates and average / standard deviation for prompts sent to the LLMs.
Figure 4: Flow similarity results obtained on the TEST and OOD sets for Outline and Outline with inputs.
4.3 Metrics
We devised our own metric called Flow Similarity (FlowSim) to compare generated and expected workflows. As workflows can be represented as trees, similar to how computer programs can be represented as abstract syntax trees, we use the tree edit distance algorithm Zhang and Shasha (1989) to compute a similarity score, where higher is better. We found that this metric correlates well with human evaluation (details in Appendix A). A labeled expected workflow is required per user requirement, but evaluation can be automated. Appendix B includes a sample workflow tree.
@@ 5 Results
Figure 5: Percentage of examples with structure errors on the TEST and OOD sets. Lower is better.
@@
Table 3: Outline results with perfect RAG (TEST set). Best results are in bold and second best are underlined.
6 Error Analysis
6.1 Approach
@@
We began with a qualitative error analysis of our current production model and based on the patterns found, identified features in the ground truth dataset that appeared to negatively impact model output. These findings were then organized into a binary matrix, where 1 indicates a feature’s presence in each ground truth sample and 0, an absence (see Figure 6). We then aligned model scores with the matrix, making it easy to identify the features that could be contributing to underperformance.
@@
Figure 6: Sample binary matrix that characterizes each ground truth sample along n feature dimensions, with model scores for each sample.
6.2 Findings
@@
Table 4: Model results on three subsets of features, Structure, Input, and Enterprise. Best results are in bold and second best are underlined.
7 Conclusion
@@ Appendix A Correlation of Flow Similarity with Human Evaluation
Table 5: Correlation between human evaluation scores and FlowSim metrics. P-value is shown in parentheses.
@@ Appendix B Example of representing a workflow as a tree
Figure 7: Workflow in Figure 1 represented as a tree.
@@ Appendix C Flow Similarity Results with Structure Validation
Figure 8: FlowSim results obtained on the TEST and OOD sets for Outline and Outline with inputs, after structure validation is applied.
@@ Appendix D TEST Dataset Features
Table 6: Features identified during the error analysis process, grouped by type.

### v2

@@ Abstract.
Large Language Models, Generative AI, Retrieval-Augmented Generation, Task Decomposition, Workflows
copyright: noneconference: Workshop on Structured Knowledge for Large Language Models (SKnowLLM) at KDD 2025; August 03–07,2025; Toronto, Canadaccs: Computing methodologies Natural language processingccs: Computing methodologies Machine learningccs: Information systems Specialized information retrievalccs: Information systems Enterprise information systems
1. Introduction
Generative AI is pushing the boundaries of AI-based software. As Large Language Models (LLMs) become more capable, they have become an integral part of the software stack. Although prompting state-of-the-art models such as GPT-4o (Hurst et al., 2024), Gemini 2.0 Flash (Gemini et al., 2023), or Claude 3.5 Sonnet (Anthropic, 2024) generally suffice for traditional tasks like question-answering or summarization (Fu et al., 2024), it is not always the case on domain-specific tasks (Trad and Chehab, 2024).
While building Flow Generation, an application that generates low-code workflows based on textual user requirements, we explored whether fine-tuning a language model was necessary to achieve the desired software quality. To reduce infrastructure needs, we only considered Small Language Models (SLMs), having less than 15 billion parameters.
Figure 1. Task consists of generating a complete workflow from a user requirement.
[…]
Task consists of generating a complete workflow from a user requirement.
@@ 1. Introduction
The simplest approach to generate low-code workflows is to use prompt engineering on an off-the-shelf LLM. This would be satisfactory if the steps were high-level descriptions and the logic elements were simple. But to obtain reasonable results in this task, we had to decompose it into sub-tasks (Wies et al., 2023) and use Retrieval-Augmented Generation (RAG) (Gao et al., 2023) at various points in the pipeline. We found out that, while prompting an LLM gives reasonable results, we can do better by fine-tuning an SLM.
Our approach was to create a small yet representative training dataset for the sub-tasks comprising Flow Generation and fine-tune Mistral-Nemo-12B-Base (Mistral, 2024). For evaluation, we labeled about a thousand workflows coming from ten domains (around 100 per domain). We also created a special dataset by inviting expert enterprise users to interact with the tool and collecting data generated from their usage. We compare our approach against an instruction fine-tuned SLM as the baseline, as well as against a variety of closed and open-source LLMs, showing the benefits of fine-tuning. As our task is domain-specific, we devised a metric called Flow Similarity (FlowSim), a version of tree edit distance that represents workflows as trees.
@@
2. Related Work
Prompting (McCann et al., 2018; Brown et al., 2020) has emerged as the predominant method for using LLMs, significantly reducing the effort required compared to fine-tuning (Sanh et al., 2021; Wei et al., 2021). However, for domain-specific applications requiring extensive prior knowledge, practitioners often encounter limitations such as constrained context windows (Peng et al., 2023), which hinder the model’s ability to capture the full scope of the task. To address this, alternative approaches such as RAG (Lewis et al., 2020) or domain-specific fine-tuning (Gururangan et al., 2020) are employed to incorporate relevant knowledge. In various industry-specific settings, specialized models have demonstrated superior performance over state-of-the-art general-domain LLMs (Wu et al., 2023; Tu et al., 2024; Colombo et al., 2024). We add to this body of work by showing that a fine-tuned SLM performs better than prompting LLMs in low-code workflow generation.
@@ 2. Related Work
Evaluation methodologies vary across tasks, with each task having its own quantitative metric for model quality but also its own qualitative error analysis approach (Dou et al., 2024; Vilar et al., 2006; Gauthier-melancon et al., 2022; Bolya et al., 2020). In this work, we outline an error analysis approach for a structured output task and leverage it to gain deeper insights into model failure modes.
3. Methodology
@@ 3.1. Task Decomposition
Figure 2. Sample JSON representation of the trigger and first component steps for the workflow in Figure 1.
[…]
Sample JSON representation of the trigger and first component steps for a workflow
3.2. Retrieval-Augmented Generation
It has been shown that RAG is necessary to reduce hallucinations when generating structured outputs (Ayala and Bechard, 2024; Bassamzadeh and Methani, 2024). In our pipeline, RAG is used to suggest the following to the LLM:
@@ 3.3. Pipeline
Fine-tuning the SLM and the retriever was done using standard approaches: supervised fine-tuning and contrastive loss training.
Figure 3. System architecture with UI, AI, and data layers.
[…]
System architecture with UI, AI, and data layers
4. Experiments
4.1. Datasets
@@
Table 1. Number of examples for training (Internal and Synthetic) and evaluation (TEST and OOD).
@@
Figure 4. Flow similarity results obtained on the TEST and OOD sets for Outline and Outline with inputs.
[…]
Flow similarity results obtained on the TEST and OOD sets for Outline and Outline with inputs.
4.2. Models
For SLMs, we considered several models such as StarCoderBase-7B (Li et al., 2023) and Llama-3.1-8B (Grattafiori et al., 2024), but we chose Mistral-Nemo-12B-Base (Mistral, 2024) because this model architecture is well-supported and optimized in our enterprise system. However, we obtained similar results with other SLMs as the key ingredient for our fine-tuning is the training data.
As the baseline, we used a model of the same architecture and size that has been instruction fine-tuned but not on data from our workflow domain: Mistral-Nemo-12B-Instruct (Mistral, 2024). This allows us to validate that domain-specific data is helpful. As for LLMs, we try to cover the current landscape by comparing against:
@@ 4.2. Models
  • Open-source: Llama-3.3-70B-Instruct (Grattafiori et al., 2024)
  • Reasoning: o3-mini (OpenAI, 2025) with medium reasoning
@@
Table 2. Number of tokens in templates and average / standard deviation for prompts sent to the LLMs.
@@
4.3. Metrics
We devised our own metric called Flow Similarity (FlowSim) to compare generated and expected workflows. As workflows can be represented as trees, similar to how computer programs can be represented as abstract syntax trees, we use the tree edit distance algorithm (Zhang and Shasha, 1989) to compute a similarity score, where higher is better. We found that this metric correlates well with human evaluation (details in Appendix A). A labeled expected workflow is required per user requirement, but evaluation can be automated. Appendix B includes a sample workflow tree.
@@ 5. Results
Figure 5. Percentage of examples with structure errors on the TEST and OOD sets. Lower is better.
[…]
Percentage of examples with structure errors on the TEST and OOD sets. Lower is better.
@@
Table 3. Outline results with perfect RAG (TEST set). Best results are in bold and second best are underlined.
@@
6. Error Analysis
6.1. Approach
@@
We began with a qualitative error analysis of our current production model and based on the patterns found, identified features in the ground truth dataset that appeared to negatively impact model output. These findings were then organized into a binary matrix, where 1 indicates a feature’s presence in each ground truth sample and 0, an absence (see Table 4). We then aligned model scores with the matrix, making it easy to identify the features that could be contributing to underperformance.
Table 4. Sample binary matrix that characterizes each ground truth sample along n feature dimensions, with model scores for each sample.
@@
6.2. Findings
@@
Table 5. Model results on three subsets of features, Structure, Input, and Enterprise. Best results are in bold and second best are underlined.
@@
7. Conclusion
@@ Appendix A Correlation of Flow Similarity with Human Evaluation
Table 6. Correlation between human evaluation scores and FlowSim metrics. P-value is shown in parentheses.
@@ Appendix B Example of representing a workflow as a tree
Figure 6. Workflow in Figure 1 represented as a tree.
[…]
Workflow represented as a tree.
@@ Appendix C Flow Similarity Results with Structure Validation
Figure 7. FlowSim results obtained on the TEST and OOD sets for Outline and Outline with inputs, after structure validation is applied.
[…]
FlowSim results obtained on the TEST and OOD sets for Outline and Outline with inputs, after structure validation is applied.
@@ Appendix D TEST Dataset Features
Table 7. Features identified during the error analysis process, grouped by type.